OpenStudy (anonymous):

y'' for y= (x^2+9)^4

OpenStudy (anonymous):

Take the first derivative using the chain rule and then take the derivative of that.

OpenStudy (anonymous):

Second derivative will require both the chain rule and the product rule.

OpenStudy (anonymous):

could you show me it worked step by step

OpenStudy (anonymous):

\[y'=4(x^2+9)^3(2x)\]

OpenStudy (anonymous):

That's the first derivative. Do you see where things might have come from?

OpenStudy (anonymous):

yes i had it all right except the 2x and thats by taking the derivative of the things inside the brackets

OpenStudy (anonymous):

Yes :) that is the chain rule.

OpenStudy (anonymous):

So now we move on to the second derivative which involves the product rule.

OpenStudy (anonymous):

Do you know how to use the product rule?

OpenStudy (anonymous):

okay i got.... 4(x^2+9)^3 (2) + (2x) 12(x^2+9)^2

OpenStudy (anonymous):

Very close. First half is correct. Second half you forgot something.

OpenStudy (anonymous):

what did i forget

OpenStudy (anonymous):

oh do the product rule to finish?

OpenStudy (anonymous):

the books has 8(x^2+9)^3 + 48x^2(x^2+9)^2 = 8(x^2+9)^2(7x^2+9)

OpenStudy (anonymous):

Well the book is correct lol. All you forgot in your was the chain rule on the second half. Should have been (2x)12(x^2+9)(2x) that 2x on the end is what you forgot.

OpenStudy (anonymous):

why is the 2x at the end as well?

OpenStudy (anonymous):

Ok so when you did the second half of the product rule you can say to yourself (2x) times the derivative of the function. Well the derivative of that function 4(x^2+9)^3 is 12(x^2+9)^(2) times the chain rule (2x).

OpenStudy (anonymous):

okay so now where do they get their answer

OpenStudy (anonymous):

got the first part, but didnt get it after the simplifying

OpenStudy (anonymous):

They probably multiplied stuff out and then it eventually got them to that answer. Truthfully, I have never seen a teacher who would not accept the first answer you have. The other answer is just tedious work.

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

Actually I just looked at their second answer and you could factor...

OpenStudy (anonymous):

okay could you help me with a ln y''?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

okay. I am horrible at ln problems but here is the original y'' for y= ln x/x^2

OpenStudy (anonymous):

Ok this is a quotient rule problem or can be a product rule if you rearrange it. Take your pick.

OpenStudy (anonymous):

SO IT COULD BE (LN X)(-X^2)

OpenStudy (anonymous):

Well could be lnx(x^-2). Normally a lot of people prefer to use the product rule when possible, but it's a little easier.

OpenStudy (anonymous):

OH OKAY

OpenStudy (anonymous):

what is ln x '

OpenStudy (anonymous):

The derivative of lnx is 1/x

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

No worries :)

OpenStudy (anonymous):

okay so y' is lnx(-2x^-3) + (x^-2)(1/x)

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

okay now im going to need some help..lol

OpenStudy (anonymous):

With simplifying?

OpenStudy (anonymous):

no just straight so i can see what to do

OpenStudy (anonymous):

I'm confused. So what do you need help on? Sorry lol

OpenStudy (anonymous):

second derivative

OpenStudy (anonymous):

Ah. Ok. Well you can use the product rule again for the second derivative, but twice. So use the product rule for the lnx(-2x^-3) and then add that with the result of the product rule of (x^-2)(x^-1). I rewrote 1/x so it's easier to see.

OpenStudy (anonymous):

okay cool

OpenStudy (anonymous):

can you show me it worked out?

OpenStudy (anonymous):

Here is the working